# Magpie: Alignment Data Synthesis from Scratch by Prompting Aligned LLMs with Nothing

### v1

@@ 5 Related Work
Instruction tuning [56] and preference tuning [5] are widely used to align the responses of LLMs with human values. Instruction tuning utilizes an instruction dataset to fine-tune LLMs, where each instruction data consists of one turn or multiple turns of instructions and desired responses. The performance of instruction tuning heavily relies on the quality of instruction data [47, 53, 67]. Preference tuning further improves responses of LLMs using reinforcement learning human feedback (RLHF) [5] or preference optimization [2, 19, 23, 44] based on a preference dataset.
Alignment Dataset Construction.
@@
We classify the existing methods of creating datasets for model alignment into two main categories: human interactions with LLMs and synthetic instruction generation. To create datasets for alignment, previous studies have collected human interactions with LLMs [14, 64, 65, 66, 26]. However, manually crafting instructions is not only time-consuming and labor-intensive, but may also incorporate toxic content [64]. Another category of approaches [53, 47, 58, 59, 55, 46] focus on prompting LLMs to generate synthetic instruction datasets, beginning with a small set of human-annotated seed instructions and expanding these through few-shot prompting. However, these methods face a diversity challenge, as few-shot prompting often results in new instructions that are too similar to the original seed questions [31]. To enhance coverage, some research [16, 31] summarizes world knowledge and employs it to generate synthetic datasets. We note that our Magpie dataset also belongs to the synthetic dataset. However, we leverage the prompt template with no requirement for seed questions or prompt engineering.
@@
Compared to the above two main categories, alignment data can also be generated by transforming existing data [54, 45, 20]. However, the constrained variety of NLP tasks in these datasets may impede the ability of tuned LLMs to generalize in real-world scenarios [31]. There are also mixture datasets (e.g., [24, 49, 38, 67]) that combine or select high-quality instruction data from various existing open-source instruction datasets to enhance coverage [24, 49] and/or improve overall performance [38, 67]. There are also data construction methods focusing on improving the reasoning and math abilities [61, 62], which can be further merged with Magpie for creating a better mixture of data for instruction tuning.
Training Data Extraction.
@@
Language models have the capability to memorize examples from their training datasets, potentially enabling malicious users to extract private information [8, 7, 9]. Pioneering work [27, 9, 41] has demonstrated that it is possible to extract private pre-training data from BERT [15], GPT-2 [43], and ChatGPT [1], respectively. Yu et al. [60] propose several tricks including adjusting sampling strategies to better extract training datasets from language models. Recently, Kassem et. al. [25] propose a black-box prompt optimization method that uses an attacker LLM to extract high levels of memorization in a victim LLM. Wang et al. [52] leverage membership inference attack (MIA) to extract fine-tuning datasets from fine-tuned language models. Bai et al. [4] extracts the training dataset of production language models via special characters (e.g., structural symbols of JSON files, and , # in emails and online posts). Different from the prior work, we aim to create publicly available alignment datasets with minimal human effort by leveraging the remarkable generation capabilities of LLMs, rather than extracting private training data from LLMs.
6 Limitations and Ethical Considerations
Limitations.
In certain scenarios, users may aim to fine-tune LLMs using domain-specific instruction data. Investigating how to configure Magpie to efficiently generate the desired domain-specific instructions (e.g., math problems) is subject to our future work. Also, there is still a gap between Magpie-tuned LLMs and official Llama-3-Instruct on datasets such as WildBench and MMLU, which suggest that we should focus on producing harder reasoning tasks and feedback learning data.
License and Legality.
The instruction datasets generated by Magpie in this paper are subject to CC BY-NC license and Meta Llama 3 Community license. While users are permitted to distribute, adapt, and further develop our method Magpie, it is the responsibility of the users to apply Magpie to LLMs in compliance with the associated license agreement. We hereby disclaim any liability for misuse of data generated by users of Magpie.
Societal Impact and Potential Harmful Consequences.
@@ 6 Limitations and Ethical Considerations
The primary objective of this paper is to develop a scalable method to synthesize instruction data to enhance the instruction-following capabilities of LLMs, and thus align them with human values. However, the data generated by Magpie may contain harmful instructions and/or responses, which may lead to unsafe behaviors if used raw in instruction tuning. Our empirical evaluations indicate that such harmful data instances constitute less than 1% of the dataset. To mitigate this risk, we develop a filtering technique in Appendix B to identify and remove these instances.
@@ 8 Acknowledgement
The research of Z. Xu, F. Jiang, L. Niu, and R. Poovendran is partially supported by the National Science Foundation (NSF) AI Institute for Agent-based Cyber Threat Intelligence and Operation (ACTION) under grant IIS 2229876. The research of Y. Choi is partially supported by the National Science Foundation (NSF) under grant DMS-2134012 (Scaling Laws of Deep Learning) and the Office of Naval Research (ONR) under grant N00014-24-1-2207 (Symbolic Knowledge Distillation of LLMs for All: Diverse Scales, Skills, and Values).
@@
Appendix A Magpie Extension
In this section, we explore the extension of Magpie. We first outline the process for constructing a multi-turn dataset (Magpie-MT). We then discuss methods for controlling instruction tasks using Magpie. Finally, we will briefly discuss how to develop a preference optimization dataset based on Magpie.
A.1 Building Multi-Turn Magpie
To construct Magpie-MT, we initially follow Steps 1 and 2 to generate the first turn of instruction and response. For subsequent turns, we append the pre-query template to the end of the full prompt from the previous round of communication. We have observed that the model may occasionally forget its role as the user, especially for the 8B model. To mitigate this, we employ a system prompt designed to control the behavior of the LLM and reinforce its awareness of the multi-round conversation context. The full prompt for building the instructions of Magpie-MT can be found in Figure 11 in Appendix F. We follow the procedure described in Step 2 of Section 2 to generate responses and yield the multi-turn instruction dataset.
A.2 Control Instruction Tasks of Magpie
In some scenarios, users may wish to fine-tune large language models (LLMs) using domain-specific instruction data, such as code or mathematical content, to enhance performance within specific domains. In this section, we introduce a lightweight and effective method to control the task category of generated instructions. Our approach involves guiding LLMs through the system prompt by specifying that it is a chatbot tailored for a particular domain and outlining the types of user queries it might encounter. We provide an example of a system prompt designed to control the generation of math-related instructions, as illustrated in Figure 12 within Appendix F.
A.3 Building Preference Optimization Dataset with Magpie
Magpie can be readily adapted to create preference datasets by integrating responses generated by the instruct model with those from the base model. Specifically, utilizing the reward difference outlined in Section 3, a preference dataset can be assembled by designating the response from the instruct model as the preferred response, and the response from the base model as the less preferred one, provided that $r^{*}-r_{base}>0$. We will soon open-source Magpie-PO, a preference optimization dataset to further align LLMs with human preferences.
Appendix B Filter Setups
@@
    Task Category: The specific category of the instructions. See Appendix C.1 for details.
@@
We provide several off-the-shelf configurations, as demonstrated in Table 4. We defer the detailed performance analysis of each filter configuration for Magpie-Pro to Appendix E.2.
@@
Appendix C More Dataset Analysis
This section provides additional dataset analysis, complementing the discussions in Section 3.
C.1 Additional Analysis on Dataset Coverage and Attributes.
@@ Appendix C More Dataset Analysis
C.2 Additional Safety Analysis
Table 5 illustrates the percentage of different unsafe categories of Magpie-Air and Magpie-Pro, as labeled by Llama-Guard-2 [48]. We have two key observations. First, the proportion of data containing potentially harmful queries is minimal, with less than 1% for both datasets. Second, the majority of unsafe responses fall into the category of specialized advice, which includes responses that may offer specialized financial, medical, or legal advice, or suggest that dangerous activities or objects are safe.
Table 5: This table shows the percentage of different unsafe categories of Magpie-Air and Magpie-Pro tagged by Llama-Guard-2 [48] model.
@@ C.2 Additional Safety Analysis
C.3 Ablation Analysis on Generation Configurations
@@
Ablation Analysis on the System Prompt. Figure 10 compares the use of system prompt compared with not using it in Step 1 of Magpie. Since the Llama-3 model does not have an official system prompt, we use the default system prompt from Vicuna [10]: A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. We observe that using a system prompt generally results in a decrease in the overall quality of instructions, and the instructions are easier. Consequently, we recommend not appending system prompts in default settings.
Appendix D Detailed Experimental Setups
D.1 Experimental Setups for Generating Magpie-Air and Magpie-Pro
As detailed in Appendix C.3, varying decoding parameters in Step 1 can significantly influence the quality, difficulty, and diversity of the generated instructions. To optimize the trade-offs among these attributes, we employ diverse decoding parameters for the generation of Magpie-Air and Magpie-Pro. Table 6 presents the configurations of Magpie-Air and Magpie-Pro, showcasing how diverse decoding parameters shape each dataset.
@@ D.1 Experimental Setups for Generating Magpie-Air and Magpie-Pro
D.2 Experimental Setups for Instruction Tuning
@@
Decoding parameters for evaluation benchmarks. For Arena-Hard [32] and WildBench [34], we follow its default setting and use greedy decoding for all settings. For AlpacaEval 2 [33] which allows the model provider to specify decoding parameters, we also employ greedy decoding in all experiments with a slightly increased repetition penalty ($RP=1.2$) to mitigate the potential repetitive outputs during the generation.
Appendix E Additional Experimental Results
E.1 Performance of Magpie-MT
@@
E.2 Ablation Analysis on Filter Designs

### v2

@@ 5 Related Work
LLM Alignment. Instruction tuning (Wei et al., 2022) and preference tuning (Bai et al., 2022) are widely used to align the responses of LLMs with human values. Instruction tuning utilizes an instruction dataset to fine-tune LLMs, where each instruction data consists of one turn or multiple turns of instructions and desired responses. The performance of instruction tuning heavily relies on the quality of instruction data (Taori et al., 2023; Wang et al., 2023; Zhou et al., 2023). Preference tuning further improves responses of LLMs using reinforcement learning human feedback (RLHF) (Bai et al., 2022) or preference optimization (Azar et al., 2024; Ethayarajh et al., 2024; Hong et al., 2024; Rafailov et al., 2023) based on a preference dataset.
@@
Alignment Dataset Construction. We classify the existing methods of creating datasets for model alignment into two main categories: human interactions with LLMs and synthetic instruction generation. To create datasets for alignment, previous studies have collected human interactions with LLMs (Databricks, 2023; Zhao et al., 2024; Zheng et al., 2024; 2023; Köpf et al., 2023). However, manually crafting instructions is not only time-consuming and labor-intensive, but may also incorporate toxic content (Zhao et al., 2024). Another category of approaches (Wang et al., 2023; Taori et al., 2023; Xu et al., 2023a; b; Wang et al., 2024c; Sun et al., 2023) focus on prompting LLMs to generate synthetic instruction datasets, beginning with a small set of human-annotated seed instructions and expanding these through few-shot prompting. However, these methods face a diversity challenge, as few-shot prompting often results in new instructions that are too similar to the original seed questions (Li et al., 2024a). To enhance coverage, some research (Ding et al., 2023; Li et al., 2024a) summarizes world knowledge and employs it to generate synthetic datasets. We note that our Magpie dataset also belongs to the synthetic dataset. However, we leverage the prompt template without any requirement for seed questions or prompt engineering.
@@
Compared to the above two main categories, alignment data can also be generated by transforming existing data (Wang et al., 2022; Sanh et al., 2022; Gandhi et al., 2024). However, the constrained variety of NLP tasks in these datasets may impede the ability of tuned LLMs to generalize in real-world scenarios (Li et al., 2024a). There are also mixture datasets (e.g., (Ivison et al., 2023; Teknium, 2023a; Liu et al., 2024b; Zhou et al., 2023)) that combine or select high-quality instruction data from various existing open-source instruction datasets to enhance coverage (Ivison et al., 2023; Teknium, 2023a) and/or improve overall performance (Liu et al., 2024b; Zhou et al., 2023). There are also data construction methods focusing on improving the reasoning and math abilities (Yue et al., 2023; 2024), which can be further merged with Magpie for creating a better mixture of data for instruction tuning.
@@
Training Data Extraction. Language models have the capability to memorize examples from their training datasets, potentially enabling malicious users to extract private information (Brown et al., 2022; Biderman et al., 2023; Carlini et al., 2021). Pioneering work (Krishna et al., 2020; Carlini et al., 2021; Nasr et al., 2023) has demonstrated that it is possible to extract private pre-training data from BERT (Devlin et al., 2018), GPT-2 (Radford et al., 2018), and ChatGPT (Achiam et al., 2023), respectively. Yu et al. (2023) propose several techniques including adjusting sampling strategies to better extract training datasets from language models. Recently, Kassem et al. (2024) propose a black-box prompt optimization method that uses an attacker LLM to extract high levels of memorization in a victim LLM. Wang et al. (2024b) leverage membership inference attack (MIA) to extract fine-tuning datasets from fine-tuned language models. Bai et al. (2024) extract the training dataset of production language models via special characters (e.g., structural symbols of JSON files, and , # in emails and online posts). Different from the prior work, we aim to create publicly available alignment datasets with minimal human effort by leveraging the remarkable generation capabilities of LLMs, rather than extracting private training data from LLMs.
6 Limitations, Discussions, and Ethical Considerations
Limitations and Discussions. Magpie-aligned LLMs demonstrate strong performance on instruction following benchmarks compared to the official Llama-3-8B-Instruct. However, we observe a performance degradation on math and reasoning benchmarks. Although we leverage the techniques described in Section 2.2 to generate specialized booster reasoning datasets, there is still a performance gap between Magpie-aligned LLMs and the official models. Enhancing the reasoning ability of Magpie-aligned models presents a promising direction for future research.
@@ 6 Limitations, Discussions, and Ethical Considerations
Societal Impact and Potential Harmful Consequences. The primary objective of this paper is to develop a scalable method to synthesize instruction data to enhance the instruction-following capabilities of LLMs, and thus align them with human values. However, the data generated by Magpie may contain harmful instructions and/or responses, which may lead to unsafe behaviors if used raw in instruction tuning. Our empirical evaluations indicate that such harmful data instances constitute less than 1% of the dataset. Our data filtering technique in Appendix C can identify and remove these instances, thus mitigating the risk.
@@ 8 Acknowledgement
The research of Z. Xu, F. Jiang, L. Niu, and R. Poovendran is partially supported by the National Science Foundation (NSF) AI Institute for Agent-based Cyber Threat Intelligence and Operation (ACTION) under grant IIS 2229876 and Air Force Office of Scientific Research (AFOSR) under grant FA9550-23-1-0208. The research of Y. Choi is partially supported by the National Science Foundation (NSF) under grant DMS-2134012 (Scaling Laws of Deep Learning) and the Office of Naval Research (ONR) under grant N00014-24-1-2207 (Symbolic Knowledge Distillation of LLMs for All: Diverse Scales, Skills, and Values).
@@
Appendix A Statistics of instruction datasets generated by Magpie compared to other instruction datasets.
Magpie can be readily deployed to state-of-the-art open-weight model families, including but not limited to Llama-3 (Meta, 2024), Llama-3.1 (Dubey et al., 2024), Qwen2 (Yang et al., 2024), Gemma-2 (Team et al., 2024), and Phi-3 (Abdin et al., 2024) model families.
In what follows, we compare datasets generated by Magpie with the above model families compared to other state-of-the-art instruction datasets. The Magpie dataset family encompasses over 11.4 million diverse and high-quality instructions and corresponding responses generated from state-of-the-art open-source models. This corpus represents the largest alignment dataset for LLMs that does not rely on human-written questions or employ complex multi-stage pipelines.
Table 4: Statistics of Magpie family compared to other instruction datasets. Tokens are counted using the tiktoken library (OpenAI, 2024).
Synthetic Alpaca (Taori et al., 2023) 52K 1 Low text-davinci-003 67.38±54.88 3.5M
Evol Instruct (Xu et al., 2023a) 143K 1 Low ChatGPT 473.33±330.13 68M
UltraChat (Ding et al., 2023) 208K 3.16 Low GhatGPT 376.58±177.81 238M
Human Dolly (Databricks, 2023) 15K 1 High ChatGPT 94.61±135.84 1.42M
ShareGPT (Zheng et al., 2023) 112K 4.79 High ChatGPT 465.38±368.37 201M
WildChat (Zhao et al., 2024) 652K 2.52 High GPT-3.5 & GPT-4 727.09±818.84 852M
LMSYS-Chat-1M (Zheng et al., 2024) 1M 2.01 High Mix 260.37±346.97 496M
Mixture Deita (Liu et al., 2024b) 9.5K 22.02 - Mix 372.78±182.97 74M
OpenHermes (Teknium, 2023a) 243K 1 - Mix 297.86±258.45 72M
Tulu V2 Mixture (Ivison et al., 2023) 326K 2.31 - Mix 411.94±447.48 285M
Magpie Llama-3-Magpie-Air 3M 1 No Llama-3-8B-Instruct 426.39±217.39 1.28B
Llama-3-Magpie-Air-MT 300K 2 No Llama-3-8B-Instruct 610.80±90.61 366M
Llama-3-Magpie-Pro 1M 1 No Llama-3-70B-Instruct 478.00±211.09 477M
Llama-3-Magpie-Pro-MT 300K 2 No Llama-3-70B-Instruct 554.53±133.64  [^3] M
Llama-3.1-Magpie-Pro 1M 1 No Llama-3.1-70B-Instruct 482.35±378.45 482M
Llama-3.1-Magpie-Pro-MT 300K 2 No Llama-3.1-70B-Instruct 552.53±325.49 331M
Qwen2-Magpie-Air 3M 1 No Qwen2-7B-Instruct 577.87±416.10 1.73B
Qwen2-Magpie-Pro 1M 1 No Qwen2-72B-Instruct 424.87±339.71 424M
Gemmma-2-Magpie-Pro 1M 1 No Gemma-2-27b-it 483.90±237.80 259M
Phi-3-Magpie-Pro 534K 1 No Phi-3-Medium-Instruct 391.38±414.32 391M
Appendix B MagpieLM
Figure 6: This figure shows the performance of MagpieLM-4B-Chat and MagpieLM-8B-Chat compared with baselines. MagpieLM significantly outperforms baselines of similar sizes.
In this section, we discuss the details of our MagpieLM. To construct the SFT and DPO datasets for aligning MagpieLM, we select 550K and 200K high-quality instructions, respectively, from the Magpie family. These instructions cover diverse categories, ensuring a comprehensive training set. We then generate corresponding responses using the Gemma-2-9b-it model (Team et al., 2024).
The benchmark performance of both models is demonstrated in Figure 6. Notably, MagpieLM significantly outperforms other baselines of similar model sizes across multiple benchmarks, including Alpaca Eval 2 (Li et al., 2023b), Arena Hard (Li et al., 2024b), and Wildbench (Lin et al., 2024).
Appendix C Filter Setups
@@
    Task Category: The specific category of the instructions. See Appendix D.1 for details.
@@
We provide several off-the-shelf configurations, as demonstrated in Table 5. We defer the detailed performance analysis of each filter configuration for Magpie-Pro to Appendix F.4.
@@
Appendix D More Dataset Analysis
This section provides additional dataset analysis, complementing the discussions in Section 3. Statistics including lengths of instructions and responses are illustrated in Figure 8.
D.1 Additional Analysis on Dataset Coverage and Attributes.
Dataset Coverage Measured by T-SNE.
Figure 8 presents the t-SNE plots of Magpie-Pro, Alpaca, Evol Instruct, and UltraChat. Each t-SNE plot is generated by randomly sampling 10,000 instructions from the associated dataset. We observe that the t-SNE plot of Magpie-Pro encompasses the area covered by the plots of Alpaca, Evol Instruct, and UltraChat. This suggests that Magpie-Pro provides a broader or more diverse range of topics, highlighting its extensive coverage across varied themes and subjects.
Figure 7: Lengths of instructions and responses in Magpie-Air/Pro.
@@ Appendix D More Dataset Analysis
D.2 Additional Safety Analysis
Table 6 illustrates the percentage of different unsafe categories of Magpie-Air and Magpie-Pro, as labeled by Llama-Guard-2 (Team, 2024). We have two key observations. First, the proportion of data containing potentially harmful queries is minimal, with less than 1% for both datasets. Second, the majority of unsafe responses fall into the category of specialized advice, which includes responses that may offer specialized financial, medical, or legal advice, or suggest that dangerous activities or objects are safe.
Table 6: This table shows the percentage of different unsafe categories of Magpie-Air and Magpie-Pro tagged by Llama-Guard-2 Team (2024) model.
@@ D.2 Additional Safety Analysis
D.3 Ablation Analysis on Generation Configurations
@@
Ablation Analysis on the System Prompt. Figure 12 compares the use of system prompt compared with not using it in Step 1 of Magpie. Since the Llama-3 model does not have an official system prompt, we use the default system prompt from Vicuna (Chiang et al., 2023): A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. We observe that using a system prompt generally results in a decrease in the overall quality of instructions, and the instructions are easier. Consequently, we recommend not appending system prompts in default settings.
D.4 Impact of Annotating Models
Figure 13: This figure compares the impact of different annotators on evaluating the instruction quality and difficulty.
We note that LLMs may occasionally favor its own response (Deutsch et al., 2022). In what follows, we conduct experiments to evaluate the impact of annotating models when labeling quality and difficulty of the Magpie-Air dataset. We used the Qwen-2-7B-Instruct model (outside the Llama-3 family) to annotate the quality and difficulty of our Magpie-Air dataset. The statistics are summarized in Figure 13.
Our findings show that even when evaluated by Qwen-2-7B-Instruct, the Magpie-Air dataset maintains high quality and difficulty, which is even higher than those originally annotated by Llama-3-8B-Instruct. This suggests that our dataset’s quality is robust across different annotators.
Appendix E Detailed Experimental Setups
E.1 Experimental Setups for Generating Magpie-Air and Magpie-Pro
As detailed in Appendix D.3, varying decoding parameters in Step 1 can significantly influence the quality, difficulty, and diversity of the generated instructions. To optimize the trade-offs among these attributes, we employ diverse decoding parameters for the generation of Magpie-Air and Magpie-Pro. Table 7 presents the configurations of Magpie-Air and Magpie-Pro, showcasing how diverse decoding parameters shape each dataset.
@@ E.1 Experimental Setups for Generating Magpie-Air and Magpie-Pro
E.2 Experimental Setups for Instruction Tuning and Preference Tuning
@@
Table 9: This table shows the hyper-parameters for direct preference optimization.
Hyper-parameter Value
Learning Rate $5\times 10^{-7}$
Number of Epochs $1$
Number of Devices $4$
Per-device Batch Size $2$
Gradient Accumulation Steps $16$
Effective Batch Size $128$
Optimizer Adamw with $\beta s=(0.9,0.999)$ and $\epsilon=10^{-8}$
Learning Rate Scheduler cosine
Warmup Ratio $10\%$
Preference Tuning Hyper-parameters. Table 9 demonstrates the detailed DPO hyper-parameters for aligning Llama-3-8B using Magpie-Air-DPO and Magpie-Pro-DPO. These experiments were conducted using Alignment Handbook [^6] https://github.com/huggingface/alignment-handbook.
Decoding parameters for evaluation benchmarks. For Arena-Hard (Li et al., 2024b) and WildBench (Lin et al., 2024), we follow its default setting and use greedy decoding for all settings. For AlpacaEval 2 (Li et al., 2023b) which allows the model provider to specify decoding parameters, we also employ greedy decoding in all experiments with a slightly increased repetition penalty ($RP=1.2$) to mitigate the potential repetitive outputs during the generation.
Appendix F Additional Experimental Results
F.1 Performance of Magpie-MT
@@
F.2 Compare Magpie and Self-Instruct using Llama-3-8B-Instruct
To compare the performance of Magpie and other synthetic dataset generation methods using the same model, we follow the official Self-Instruct (Wang et al., 2023) setup and generate a 100K supervised fine-tuning dataset using Llama-3-8B-Instruct. For a fair comparison, we select the first 100K data samples from the Magpie-Air dataset generated by Llama-3-8B-Instruct. The performance of models fine-tuned with these two datasets is shown in the table 11.
Table 11: This table compares the performance of models fine-tuned using 100K instruction-following datasets generated by Self-Instruct and Magpie. All models are supervised-fine-tuned on the Llama-8B base models. We observe that Magpie significantly outperforms Self-Instruct across all benchmarks.
Dataset #Convs AlpacaEval 2 Arena-Hard
GPT-4-Turbo (1106) Llama-3-8B-Instruct
LC (%) WR (%) SD LC (%) WR (%) SD WR (%)
Magpie-Air-100K 100K 20.17 21.33 1.21 46.82 48.76 1.44 15.7
Self-Instruct (Llama-3) 100K 7.21 5.18 0.7 17.86 12.73 1.05 4.0
We observe a significant performance gap between models fine-tuned with datasets generated by Self-Instruct and our Magpie. Our analysis revealed that the instruction format in Self-Instruct-generated datasets is predominantly constrained by the patterns defined in the seed instructions, resulting in a lack of diversity. This comparison indicates the novelty of our Magpie in generating diverse high-quality instructions without any seed questions.
F.3 Ablation Analysis on Data Quantity and Quality
In what follows, we compare within the family of datasets generated by Magpie in Table 12. These datasets differ in size, deployment of filtering, and models used to generate data. We observe that as the dataset’s size increases, the fine-tuned model’s performance improves, indicating that data quantity plays a critical role in enhancing instruction-following capabilities. Furthermore, the model fine-tuned with Magpie-Pro-300K-Filtered outperforms those fine-tuned with the same or even higher amounts of raw data. This demonstrates the effectiveness of our filtering technique, and underscores the importance of data quality. Finally, we observe that the models fine-tuned with Magpie-Pro consistently outperform those fine-tuned with Magpie-Air. The reason is that Magpie-Pro is generated by the more capable model, i.e., Llama-3-70B-Instruct.
Table 12: This table compares Magpie datasets within its family that differ in size, deployment of filtering, and models used to generate data. All models are supervised-fine-tuned on the Llama-8B base models.
Dataset #Convs AlpacaEval 2 Arena-Hard
GPT-4-Turbo (1106) Llama-3-8B-Instruct
LC (%) WR (%) SD LC (%) WR (%) SD WR(%)
Magpie-Air 300K-Raw 300K 21.99 21.65 1.21 48.63 48.06 1.42 15.8
3M-Raw 3M 22.96 21.09 1.20 50.57 48.40 1.42 16.1
300K-Filtered 300K 22.66 23.99 1.24 49.27 50.8 1.44 14.9
Magpie-Pro 300K-Raw 300K 21.65 22.19 1.2 49.65 50.84 1.42 15.9
1M-Raw 1M 24.16 23.93 1.25 49.97 50.34 1.43 16.7
100K-Filtered 100K 20.47 24.52 1.25 47.92 52.75 1.43 17.2
200K-Filtered 200K 22.11 26.02 1.26 51.17 56.76 1.41 15.9
300K-Filtered 300K 25.08 29.47 1.35 52.12 53.43 1.44 18.9
Magpie-Air + Magpie-Pro 4M-Raw 4M 24.45 24.08 1.26 51.96 52.08 1.42 15.5
F.4 Ablation Analysis on Filter Designs